OpenStudy (anonymous):

Solve. 1/2^x = 1/(x+2)

OpenStudy (anonymous):

cross multiply x+2 = 2^x x = 2

OpenStudy (anonymous):

how did you get from x + 2 = 2^x to x = 2?

OpenStudy (anonymous):

think about it, what exponent does x need to be to make it work in that equation. Only 2 works :)

OpenStudy (anonymous):

is there some mathematical way to prove that so you dont lose marks for communication. Canadian problems --.-

OpenStudy (anonymous):

not with algebra, graphing yes

OpenStudy (anonymous):

oh ok thanks man :)